OpenStudy (anonymous):

Which of the following numbers is prime? A. 133 B. 143 C. 153 D. 163 C?

OpenStudy (mathstudent55):

Divide each number by 2, 3, 4, 5, 6, ...

OpenStudy (mathstudent55):

It's not C.

OpenStudy (anonymous):

which number can you not divide evenly= correct answer

OpenStudy (mathstudent55):

153/3 = 51

OpenStudy (anonymous):

Why 153 is not a prime number? Because 153 has divisors rather than 1 and itself. 153 can be divided by 3 and 17. what i found

OpenStudy (anonymous):

It's D, (163)

OpenStudy (anonymous):

its 163

OpenStudy (mathstudent55):

correct

OpenStudy (anonymous):

so wait its 163?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yea

OpenStudy (anonymous):

I meant to say a different letter sorry. But if you look at it if I am correct its D) 163.

OpenStudy (anonymous):

Because 163 has no positive divisors rather than 1 and itself

OpenStudy (mathstudent55):

133/7 = 19 143/11 = 13 153/3 = 17 Answer is 163

OpenStudy (anonymous):

cool guys thank! make sure to give each other medals and stuff cause you are all awesome ! i will make sure to fan you all!

OpenStudy (mathstudent55):

You're welcome.

OpenStudy (mathstudent55):

To find if a number is prime, you need to test it by dividing it by all integers up to the square root of the number. For some of the integers that you need to divide by, there are tests of divisibility that will save you time. For example, let's test 143 to see if it's prime. sqrt(143) = 11.958... That means that for 143, you only need to test up to integer 11. Division by 1 is unnecessary. Since 143 is odd, it's not divisible by 2, 4, 6, 8, 10, so all those integers don't need to be tested. For divisibility by 3 there is a test. 143 does not pass it, so it's not divisible by 3. Automatically, it's not divisible by 9 either. For 5 there is a test, and 143 is not divisible by 5. There is no test for divisibility by 7, so you must try it. 143/7 = 20.42857..., so it's not divisible by 7. That only leaves 11. 143/11 = 13 143 is divisible by 11. This shows it's not prime.